# Efficient Large-Scale Urban Parking Prediction: Graph Coarsening Based on Real-Time Parking Service Capability

### v1

@@ Abstract
With the sharp increase in the number of vehicles, the issue of parking difficulties has emerged as an urgent challenge that many cities need to address promptly. In the task of predicting large-scale urban parking data, existing research often lacks effective deep learning models and strategies. To tackle this challenge, this paper proposes an innovative framework for predicting large-scale urban parking graphs leveraging real-time service capabilities, aimed at improving the accuracy and efficiency of parking predictions. Specifically, we introduce a graph attention mechanism that assesses the real-time service capabilities of parking lots to construct a dynamic parking graph that accurately reflects real preferences in parking behavior. To effectively handle large-scale parking data, this study combines graph coarsening techniques with temporal convolutional autoencoders to achieve unified dimension reduction of the complex urban parking graph structure and features. Subsequently, we use a spatio-temporal graph convolutional model to make predictions based on the coarsened graph, and a pre-trained autoencoder-decoder module restores the predicted results to their original data dimensions, completing the task. Our methodology has been rigorously tested on a real dataset from parking lots in Shenzhen. The experimental results indicate that compared to traditional parking prediction models, our framework achieves improvements of 46.8% and 30.5% in accuracy and efficiency, respectively. Remarkably, with the expansion of the graph’s scale, our framework’s advantages become even more apparent, showcasing its substantial potential for solving complex urban parking dilemmas in practical scenarios.
@@ I INTRODUCTION
To alleviate the parking difficulty issue, urban-level smart parking platforms [2] have emerged. By leveraging cutting-edge technologies such as digital twins, the Internet of Things, and cloud computing [3], these platforms have successfully integrated urban parking facility resources, effectively solving the problem of information asymmetry between users and available parking spaces. Taking the Wuhan smart parking project [4] as an example, by the end of 2022, the platform had integrated more than 4,000 parking lots, including approximately 480,000 parking space data. With its portability, the platform offers convenience to citizens, who can easily check real-time parking availability in different areas through their mobile phones and use services such as parking reservation [5] and parking navigation [6],[7],[8] to achieve precise parking. In these platforms, the key to providing smart parking services lies in parking prediction technology [9]. Indeed, parking prediction has always been a hot research topic in urban parking technology [10],[11]. By effectively predicting the availability of parking spaces in future time periods, it can not only enhance user experience but also improve parking turnover, optimize drivers’ parking behavior, and thereby help to allocate urban parking resources rationally.
Recent studies have framed such parking prediction tasks as spatio-temporal graph modeling problems [12]. On one hand, this is because parking lots are connected through urban road networks, forming unique parking graphs [13], for which graph convolutional networks can be utilized to capture the spatial characteristics of the parking graph [14]. On the other hand, given the tidal characteristics of parking flow, recurrent neural networks can be employed to fit the time series regression of parking data [15]. This approach, which considers both spatial and temporal dependencies, is known as spatio-temporal Graph Convolutional Networks (GCNs).
However, for the vehicle-dense cities mentioned above, with the rapid development of wireless sensors and the widespread use of vehicles [16], a massive influx of real-time parking data continuously flows into urban-level parking platforms. This leads to a dramatic increase in the volume of parking data managed by existing platforms and the coverage range of the parking Internet of Things network [17]. The rapid influx of such information necessitates an unsustainable amount of time to train spatio-temporal GCNs at the required scale for large-scale urban parking prediction.
@@
  • We have proposed a method for constructing parking graphs that leverages a ParkingRank graph attention mechanism. This method intricately integrates the real-time service capacity assessment of parking lots into a graph attention network, creating a parking graph that accurately mirrors real-world parking behavior preferences. This graph is adept at capturing the complexity of parking scenarios, while also remaining flexible enough to adapt to its dynamic shifts, laying the foundation for downstream coarsening as well as prediction tasks.
@@
This paper is structured as follows: Section II provides a concise overview of spatio-temporal GCNs for parking prediction, techniques for dimensionality reduction in large-scale graphs, and the applications of Autoencoders (AE). Section III delves into the specifics of our innovative ParkingRank graph attention model and outlines our parking prediction methodology, which leverages TCN-AE for coarsening. Section IV details the experimental framework and analysis employed to assess the effectiveness of our proposed approach. We conclude by summarizing the key insights and contributions of this study.
@@ II RELATED WORK
II-A Spatio-temporal Graph Convolutional Models
Spatio-temporal GCNs [31] have emerged as a powerful and prevalent approach within the realm of traffic prediction, adept at navigating the intricacies of traffic graphs by integrating both spatial and temporal dependencies. Unlike relying solely on GCNs or RNNs, spatio-temporal GCNs excel at dissecting time-series data embedded in graph structures [32], thereby yielding more precise prediction. For instance, Zhao et al. [33] introduced the T-GCN model, ingeniously combining GCN with GRU to understand the intricate topology of traffic networks and the dynamic shifts in traffic data, enhancing the capture of spatio-temporal characteristics more effectively. In a similar vein, Sun et al. [34] unveiled DDSTGCN, leveraging TCN to track variations in traffic states alongside GCN for topology learning, markedly boosting traffic prediction accuracy. In recent years, scholars have suggested employing attention mechanisms to bolster the extraction of spatio-temporal features, thereby extending the model’s interpretability. Zheng et al. [35] proposed a novel strategy that incorporates spatial and temporal attention layers to elucidate the spatio-temporal correlations within heterogeneous traffic flows, thereby not only elevating predictive performance but also the interpretability of the findings. Furthermore, Li et al. [36] developed AST-GAT, incorporating multi-head graph attention and attention-based LSTM modules to intricately map the spatio-temporal interdependencies among road segments, aiming to refine the precision of traffic prediction.
Due to the inherent structural and functional similarities between urban transportation networks and urban parking networks, both involve complex spatial relationships and dynamically changing temporal characteristics. Leveraging advanced spatio-temporal graph convolutional models from the field of traffic prediction, we believe can effectively enhance the accuracy of parking prediction. However, in the domain of parking prediction, the industrial sector continues to prioritize concerns regarding time costs and resource consumption, with a current lack of effective solutions. These challenges will limit the feasibility and effectiveness of the aforementioned spatio-temporal graph convolutional models in practical industrial applications.
@@ II-C Applications of Autoencoders
AEs have demonstrated remarkable proficiency in managing intricate spatio-temporal data relevant to transportation systems [35]. Typically employing symmetrically structured encoders and decoders for unsupervised learning, these AEs are adept not only at distilling input data into compact representations of potential vectors [27], but also at restoring the data to its original dimensions. They achieve this by optimizing the reconstruction of the data’s objective, thereby minimizing loss [29]. For instance, the TGAE introduced by Wang et al. [43] adeptly captures the underlying patterns of traffic flow through a spatio-temporal AE, facilitating the prediction of future traffic information. Similarly, the TCN-AE devised by Mo et al. [44] leverages TCNs within its encoding module to extract features from temporal data, subsequently utilizing the decoding module to map latent representations back to the original data space, thus enabling anomaly detection.
Inspired by the capabilities of AEs in dimensionality reduction and key feature extraction of spatio-temporal data, we propose the integration of AEs into large-scale urban parking prediction tasks, aiming to address the limitations of traditional coarsening methods in handling parking lot node features. Additionally, by leveraging the advantages of TCN in processing complex temporal data, AEs can more accurately capture and analyze the parking patterns within different parking lots, thereby significantly enhancing the performance and depth of understanding of downstream parking prediction models.
@@ III-A Problem Definition
III-B ParkingRank Graph Attention
GAT [45] is an advanced spatial-based graph neural network methodology, whose primary advantage lies in redefining the aggregation of node features. Unlike GCN that typically assign equal weights to all neighboring nodes for information aggregation, GAT introduces an attention mechanism. This allows the model to dynamically allocate varying weights based on the importance of each neighboring node. Such a strategy enables GAT to more effectively capture spatial correlations within road networks and selectively emphasize those nodes that are crucial for the current task. In the application of urban parking network prediction, this feature of GAT is particularly significant. It enables the network to flexibly capture important relationships between parking nodes, focusing on those nodes that are critical for the current prediction task. We believe that this not only helps in more accurately describing the correlations between different parking lots but also aids in constructing an efficient parking network that reflects the relevancy of parking spaces. The overall framework of our proposed PRGAT is shown in Fig. 1:
Figure 1: The structure of PRGAT.
@@ III-B ParkingRank Graph Attention
The ParkingRank model [25] is an algorithm to quantitatively assess the service capability of different parking lots. By integrating complex parking information such as the total number of parking spaces in the parking lot, the degree of openness to the public, and the price of parking, it can help drivers to understand the actual situation of the parking lot in a more comprehensive way, so that they can make informed choices in the parking decision-making process. Specifically, the algorithm primarily focuses on three aspects to describe the real-time service capacity of parking lots, as shown in Equation 2:
@@
We believe that the process of quantifying the service capability of a parking lot is essentially a process of extracting the redundant parts of the parking lot features $x_{t}$, and the low-dimensional evaluation results can be used as new abstract features to replace the original high-dimensional parking data $x_{t}$.
Moreover, while GAT typically utilize Multi-Layer Perceptrons (MLP) or cosine similarity to ascertain the degree of association between nodes, such conventional methods may fall short in terms of interpretability, particularly when applied to entities like parking lots with distinct attributes. These methods often fail to offer a clear insight into the dynamics of parking lot interactions. To mitigate the black-box issue encountered by GATs in analyzing parking graphs, we integrate the parking spatio-temporal transfer matrix [2] into the computation of attention coefficients within GATs. This matrix meticulously accounts for the spatio-temporal evolution of parking lot features, encompassing real-time occupancy rates, service capabilities, and spatial connections, thereby facilitating a dynamic representation of parking cruising behavior. This nuanced incorporation allows for a richer, more detailed understanding of parking dynamics, effectively translating the raw data into actionable insights. The formulation of this transfer matrix $A_{trf}$ is presented in Equation 3.
@@
Therefore, in this paper, the input to PRGAT consists of the original adjacency matrix $A$ based on Euclidean distance and the new parking data $X^{\text{low}}\in\mathbb{R}^{T\times N\times F^{\text{low}}}$, which is composed of the service capacity $PR_{i}$, the real-time occupancy rate $q_{i}$, and latitude and longitude coordinates. The output is the updated parking data $X^{\text{re}}\in\mathbb{R}^{T\times N\times F^{\text{re}}}$. Here, $F^{\text{low}}$ represents the new feature dimensions after quantification and extraction, while $F^{\text{re}}$ denotes the new feature dimensions relearned by PRGAT.
@@
We input both the parking spatio-temporal transfer matrix $A_{\text{trf}}$ and the aforementioned attention matrix $A_{\text{attn}}$ into the softmax activation function simultaneously to derive the ParkingRank attention matrix $A_{\text{combined}}\in\mathbb{R}^{N\times N}$, as indicated in Equation 6.
@@
Algorithm 1 ParkingRank Graph Attention (PRGAT)
@@
6:     $PR_{i}\leftarrow\text{ParkingServiceCapacity}(X(i))$
7:     ${X}^{\text{low}}(i)\leftarrow\text{ParkingReconstruction}(PR_{i},X(i))$
8:  end for
9:  for $i\leftarrow 1\text{ to }N$ do
10:     $q_{i}\leftarrow\text{ParkingOccupancyRate}({X}^{\text{low}}(i))$
11:     for $j\leftarrow 1\text{ to }N$ do
12:        $s_{i,j}\leftarrow\text{Normalized}(A(i,j))$
13:        $\alpha_{i,j}^{\text{trf}}\leftarrow s_{i,j}\ast(1-q_{i})\ast PR_{i}$
14:        $e_{i,j}\leftarrow\text{Concat}(W{X}^{\text{low}}(i),W{X}^{\text{low}}(j))$
15:        $\alpha_{i,j}^{\text{ReLU}}\leftarrow\text{LeakyReLU}(a^{T}\ast e_{i,j})$
16:        $\alpha_{i,j}^{\text{Masked}}\leftarrow\text{MaskedFill}(\alpha_{i,j}^{\text{ReLU}},A(i,j))$
17:        $\alpha_{i,j}^{\text{attn}}\leftarrow\text{Normalized}(\alpha_{i,j}^{\text{Masked}})$
18:        $A_{\text{combined}}(i,j)\leftarrow\text{softmax}(\alpha_{i,j}^{\text{attn}}+\alpha_{i,j}^{\text{trf}})$
19:     end for
20:  end for
21:  ${X}^{\text{re}}\leftarrow\sigma(A_{\text{combined}}W{X}^{\text{low}})$
22:  return $A_{\text{combined}},{X}^{\text{re}}$
III-C Parking graph Coarsening

### v2

@@ Abstract
With the sharp increase in the number of vehicles, the issue of parking difficulties has emerged as an urgent challenge that many cities need to address promptly. In the task of predicting large-scale urban parking data, existing research often lacks effective deep learning models and strategies. To tackle this challenge, this paper proposes an innovative framework for predicting large-scale urban parking graphs leveraging real-time service capabilities, aimed at improving the accuracy and efficiency of parking predictions. Specifically, we introduce a graph attention mechanism that assesses the real-time service capabilities of parking lots to construct a dynamic parking graph that accurately reflects real preferences in parking behavior. To effectively handle large-scale parking data, this study combines graph coarsening techniques with temporal convolutional autoencoders to achieve unified dimension reduction of the complex urban parking graph structure and features. Subsequently, we use a spatiotemporal graph convolutional model to make predictions based on the coarsened graph, and a pre-trained autoencoder-decoder module restores the predicted results to their original data dimensions, completing the task. Our methodology has been rigorously tested on a real dataset from parking lots in Shenzhen. The experimental results indicate that compared to traditional parking prediction models, our framework achieves improvements of 46.8% and 30.5% in accuracy and efficiency, respectively. Remarkably, with the expansion of the graph’s scale, our framework’s advantages become even more apparent, showcasing its substantial potential for solving complex urban parking dilemmas in practical scenarios.
@@ I INTRODUCTION
To alleviate the parking difficulty issue, urban-level smart parking platforms [2] have emerged. By leveraging cutting-edge technologies such as digital twins, the Internet of Things, and cloud computing [3], these platforms have successfully integrated urban parking facility resources, effectively solving the problem of information asymmetry between users and available parking spaces. Taking the Wuhan smart parking project [4] as an example, by the end of 2022, the platform had integrated more than 4,000 parking lots, including approximately 480,000 parking space data. With its portability, the platform offers convenience to citizens, who can easily check real-time parking availability in different areas through their mobile phones and use services such as parking reservation [5] and parking navigation [6],[7],[8] to achieve precise parking. In these platforms, the key to providing smart parking services lies in parking prediction technology [9]. Indeed, parking prediction has always been a hot research topic in urban parking technology [10]. By effectively predicting the availability of parking spaces in future time periods, it can not only enhance user experience but also improve parking turnover, optimize drivers’ parking behavior, and thereby help to allocate urban parking resources rationally.
Recent studies have framed such parking prediction tasks as spatiotemporal graph modeling problems [11]. On one hand, this is because parking lots are connected through urban road networks, forming unique parking graphs [12], for which graph convolutional networks can be utilized to capture the spatial characteristics of the parking graph [13]. On the other hand, given the tidal characteristics of parking flow, recurrent neural networks can be employed to fit the time series regression of parking data [14]. This approach, which considers both spatial and temporal dependencies, is known as spatiotemporal Graph Convolutional Networks (GCNs).
However, for the vehicle-dense cities mentioned above, with the rapid development of wireless sensors and the widespread use of vehicles [15], a massive influx of real-time parking data continuously flows into urban-level parking platforms. This leads to a dramatic increase in the volume of parking data managed by existing platforms and the coverage range of the parking Internet of Things network [16]. The rapid influx of such information necessitates an unsustainable amount of time to train spatiotemporal GCNs at the required scale for large-scale urban parking prediction.
@@
  • We have proposed a method for constructing parking graphs that leverages a ParkingRank-based attention mechanism. Specifically, this method integrates real-time assessments of parking lot service capacities into a graph attention network, generating a parking graph that accurately reflects real-world parking behavior preferences. This graph not only captures the complex characteristics of parking scenarios but also enhances the interpretability of the graph attention mechanism in analyzing intricate parking behavior patterns. Additionally, it adapts seamlessly to dynamic changes within parking environments, thereby providing a solid foundation for downstream parking graph coarsening as well as prediction tasks.
@@
This paper is structured as follows: Section II provides a concise overview of spatiotemporal GCNs for parking prediction, techniques for dimensionality reduction in large-scale graphs, and the applications of Autoencoders (AE). Section III delves into the specifics of our innovative ParkingRank graph attention model and outlines our parking prediction methodology, which leverages TCN-AE for coarsening. Section IV details the experimental framework and analysis employed to assess the effectiveness of our proposed approach. We conclude by summarizing the key insights and contributions of this study.
@@ II RELATED WORK
II-A spatiotemporal Graph Convolutional Models
spatiotemporal GCNs [30] have emerged as a powerful and prevalent approach within the realm of traffic prediction, adept at navigating the intricacies of traffic graphs by integrating both spatial and temporal dependencies. Unlike relying solely on GCNs or RNNs, spatiotemporal GCNs excel at dissecting time-series data embedded in graph structures [31], thereby yielding more precise prediction. For instance, Zhao et al. [32] introduced the T-GCN model, ingeniously combining GCN with GRU to understand the intricate topology of traffic networks and the dynamic shifts in traffic data, enhancing the capture of spatiotemporal characteristics more effectively. In a similar vein, Sun et al. [33] unveiled DDSTGCN, leveraging TCN to track variations in traffic states alongside GCN for topology learning, markedly boosting traffic prediction accuracy. In recent years, scholars have suggested employing attention mechanisms to bolster the extraction of spatiotemporal features, thereby extending the model’s interpretability. Furthermore, Li et al. [34] developed AST-GAT, incorporating multi-head graph attention and attention-based LSTM modules to intricately map the spatiotemporal interdependencies among road segments, aiming to refine the precision of traffic prediction.
Due to the inherent structural and functional similarities between urban transportation networks and urban parking networks, both involve complex spatial relationships and dynamically changing temporal characteristics. Leveraging advanced spatiotemporal graph convolutional models from the field of traffic prediction, we believe can effectively enhance the accuracy of parking prediction. However, in the domain of parking prediction, the industrial sector continues to prioritize concerns regarding time costs and resource consumption, with a current lack of effective solutions. These challenges will limit the feasibility and effectiveness of the aforementioned spatiotemporal graph convolutional models in practical industrial applications.
@@ II-C Applications of Autoencoders
AEs have demonstrated remarkable proficiency in managing intricate spatiotemporal data relevant to transportation systems [41]. Typically employing symmetrically structured encoders and decoders for unsupervised learning, these AEs are adept not only at distilling input data into compact representations of potential vectors [26], but also at restoring the data to its original dimensions. They achieve this by optimizing the reconstruction of the data’s objective, thereby minimizing loss [28]. For instance, the TGAE introduced by Wang et al. [41] adeptly captures the underlying patterns of traffic flow through a spatiotemporal AE, facilitating the prediction of future traffic information. Similarly, the TCN-AE devised by Mo et al. [42] leverages TCNs within its encoding module to extract features from temporal data, subsequently utilizing the decoding module to map latent representations back to the original data space, thus enabling anomaly detection.
Inspired by the capabilities of AEs in dimensionality reduction and key feature extraction of spatiotemporal data, we propose the integration of AEs into large-scale urban parking prediction tasks, aiming to address the limitations of traditional coarsening methods in handling parking lot node features. Additionally, by leveraging the advantages of TCN in processing complex temporal data, AEs can more accurately capture and analyze the parking patterns within different parking lots, thereby significantly enhancing the performance and depth of understanding of downstream parking prediction models.
@@ III-A Problem Definition
III-B Parking Service Capacity Assessment
@@
The ParkingRank algorithm [24] is an algorithm to quantitatively assess the service capability of different parking lots. By integrating complex parking information such as the total number of parking spaces in the parking lot, the degree of openness to the public, and the price of parking, it can help drivers to understand the actual situation of the parking lot in a more comprehensive way, so that they can make informed choices in the parking decision-making process. Specifically, the algorithm primarily focuses on three aspects to describe the real-time service capacity of parking lots, as shown in Equation 2:
@@
We believe that the process of quantifying the service capability of a parking lot is essentially a process of extracting the redundant parts of the parking lot features $x_{t}$, and the low-dimensional evaluation results can be used as new abstract features to replace the original high-dimensional parking data $x_{t}$, providing a more concise and information-rich feature input for subsequent graph modeling.
III-C ParkingRank Graph Attention Networks
GAT [43] are an advanced spatial-based graph neural network methodology, with the core advantage of redefining the aggregation of node features. Unlike GCN, GAT introduces an attention mechanism, allowing the model to dynamically allocate different weights based on the importance of each neighboring node. In contrast, GCN typically assigns equal weights to all neighboring nodes for information aggregation. This strategy enables GAT to more effectively capture spatial correlations within road networks and selectively emphasize nodes that are more critical to the current task. This feature of GAT is particularly important in the application of urban parking network prediction. However, since GAT often uses MLP or cosine similarity to calculate the degree of association between nodes, conventional methods lack interpretability when applied to nodes such as parking lots, which have specific attributes. These methods often fail to provide an intuitive understanding of the interactions between parking lots.
To mitigate the black-box issue encountered by GATs in analyzing parking graphs, we integrate the parking spatiotemporal transfer matrix [2] into the computation of attention coefficients within GATs. This matrix meticulously accounts for the spatiotemporal evolution of parking lot features, encompassing real-time occupancy rates, service capabilities, and spatial connections, thereby facilitating a dynamic representation of parking cruising behavior. This nuanced incorporation allows for a richer, more detailed understanding of parking dynamics, effectively translating the raw data into actionable insights. The formulation of this transfer matrix $A_{trf}$ is presented in Equation 3.
@@ III-C ParkingRank Graph Attention Networks
By introducing the parking spatiotemporal transfer matrix, the GAT model can dynamically consider multiple factors, including the spatial relationships between parking lots, real-time occupancy rates, service capacities, and other relevant factors, when assigning attention weights. This mechanism enables the model to more accurately capture the spatiotemporal dependencies of nodes in complex parking networks and respond in real-time to fluctuations in parking demand. Through this approach, the model not only identifies the most critical nodes for the current prediction task but also flexibly adjusts its focus on different parking lots, optimizing the model’s representation of the parking network’s complexity and enhancing its performance in downstream prediction tasks. We believe that the improved graph attention mechanism will help construct a parking network that fully reflects the complex spatiotemporal correlations of parking lots. The overall framework of our proposed PRGAT is shown in Fig. 1:
Figure 1: The structure of PRGAT.
In this paper, the input to PRGAT consists of the original adjacency matrix $A$ based on Euclidean distance and the new parking data $X^{\text{low}}\in\mathbb{R}^{T\times N\times F^{\text{low}}}$, which is composed of the service capacity $PR_{i}$, the real-time occupancy rate $q_{i}$, and latitude and longitude coordinates. The output is the updated parking data $X^{\text{re}}\in\mathbb{R}^{T\times N\times F^{\text{re}}}$. Here, $F^{\text{low}}$ represents the new feature dimensions after quantification and extraction, while $F^{\text{re}}$ denotes the new feature dimensions relearned by PRGAT.
@@
We input both the parking spatiotemporal transfer matrix $A_{\text{trf}}$ and the aforementioned attention matrix $A_{\text{attn}}$ into the softmax activation function simultaneously to derive the ParkingRank attention matrix $A_{\text{combined}}\in\mathbb{R}^{N\times N}$, as indicated in Equation 6.
@@
Algorithm 1 ParkingRank Graph Attention Networks (PRGAT)
@@
[…]
7:     $PR_{i}\leftarrow\exp\left(s_{i}\right)\frac{y_{i}/\|y\|}{1+z_{i}/\|z\|}$
8:     $X_{i}^{\text{low}}\leftarrow\text{Concat}\left(X_{i}[:,\{\text{all}\}\backslash\{\text{service range},\right.$
[…]
10:  end for
11:  for $i\leftarrow 1\text{ to }N$ do
12:     $q_{i}\leftarrow\frac{X_{i}[:,\{\text{ occupied parking spaces }\}]}{X_{i}[:,\{\text{ parking capacity }\}]}$
13:     for $j\leftarrow 1\text{ to }N$ do
14:        $w_{i,j}\leftarrow\text{Normalized}(A(i,j))$
15:        $\alpha_{i,j}^{\text{trf}}\leftarrow w_{i,j}\ast(1-q_{i})\ast PR_{i}$
16:        $e_{i,j}\leftarrow\text{Concat}(W{X}^{\text{low}}(i),W{X}^{\text{low}}(j))$
17:        $\alpha_{i,j}^{\text{ReLU}}\leftarrow\text{LeakyReLU}(a^{T}\ast e_{i,j})$
18:        $\alpha_{i,j}^{\text{Masked}}\leftarrow\text{MaskedFill}(\alpha_{i,j}^{\text{ReLU}},A(i,j))$
19:        $\alpha_{i,j}^{\text{attn}}\leftarrow\text{Normalized}(\alpha_{i,j}^{\text{Masked}})$
20:        $A_{\text{combined}}(i,j)\leftarrow\text{softmax}(\alpha_{i,j}^{\text{attn}}+\alpha_{i,j}^{\text{trf}})$
21:     end for
22:  end for
23:  ${X}^{\text{re}}\leftarrow\sigma(A_{\text{combined}}W{X}^{\text{low}})$
24:  return $A_{\text{combined}},{X}^{\text{re}}$
III-D Parking Graph Coarsening